A small mailbag is released from a helicopter that is descending steadily at 1.23 m/s.

(a) After 4.00 s, what is the speed of the mailbag?  

(b) How far is it below the helicopter?

(c) What are your answers to parts (a) and (b) if the helicopter is rising steadily at 1.23 m/s?
Answer #1

Solution)

a) We know,

v = v0 + a*t = 1.23 + 10*4 = 41.23 m/s

b) s = a*t²/2 = 10*4^2/2 = 80 m

c) here,

s = v0*t + a*t²/2 = 1.23*4 + 10*4^2/2 = 84.92 m

=========

Good luck!:)